# Optimal Stopping with a Predicted Prior

### v1

@@
Optimal Stopping with a Predicted Prior
@@ 1 Introduction
Optimal stopping problems study a fundamental dilemma in decision‑making: to seize an expiring opportunity now, or to wait for a better chance that may never come. A classic example is the secretary problem. In this problem, $n$ values are revealed to a decision maker in a random order. After observing each value, the decision maker immediately decides whether to accept the value. The decision maker can accept at most one value and aims to maximize its expectation. An optimal solution is Dynkin’s algorithm [dynkinOptimumChoiceInstant1963, fergusonWhoSolvedSecretary1989], which rejects the first $\lfloor\frac{n}{\mathrm{e}}\rfloor$ values, and then accepts the first value that exceeds all previous ones. It guarantees that the expected accepted value is at least $\frac{1}{\mathrm{e}}$ times the maximum. This ratio of the algorithm’s expected accepted value to the maximum value in hindsight is called the competitive ratio.
In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from known prior distributions. In the i.i.d. case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [hillComparisonsStopRule1982, kertzStopRuleSupremum1986, correaPostedPriceMechanisms2017, liuVariableDecompositionProphet2021], substantially better than prior-free secretary guarantee of $\frac{1}{\mathrm{e}}\approx 0.368$.
On the one hand, the secretary problem does not capture the abundant data available to decision makers in this era. On the other hand, the assumption of complete knowledge of the prior in prophet inequality is idealized. What real-world decision makers often have are machine-learned and potentially erroneous predicted priors. duttingPostedPricingProphet2019 explored this direction for the general prophet inequality problem. They examined how mispredicted priors affect existing methods and showed a degradation in performance proportional to the prediction error. This can be viewed as an analysis of the smoothness of algorithms under the framework of algorithms with predictions [mitzenmacher2022algorithms].
@@ 1.1 Optimal Stopping with a Predicted Prior
The $\frac{1}{\mathrm{e}}$-competitive Dynkin’s algorithm is $\frac{1}{\mathrm{e}}$-consistent and $\frac{1}{\mathrm{e}}$-robust for both objectives. The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by gilbertRecognizingMaximumSequence are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent respectively, but only 0-robust (see Section 2). Further, randomly running either Dynkin’s algorithm or an i.i.d. prophet inequality algorithm linearly interpolates between their consistency and robustness. Finally, the optimality of these competitive ratios implies that no algorithm is better than $\frac{1}{\mathrm{e}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb. These baseline algorithms and impossibilities leave a significant gap between them.
@@ 1.2 Overview of Results
We define a family of bi-criteria algorithms parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Upon observing each value and when no value has been accepted, a bi-criteria algorithm accepts the current value if:
@@
    It is best-so-far, i.e., greater than or equal to all previous values; (prior-free criterion)
@@
For now, readers may interpret $t$ as the time step by which the value is observed. Later, we will also consider a continuous-time model to simplify the analysis (see Section 2 for a discussion).
We set the thresholds as follows. Before some time $\lambda_{1}$, the threshold is $1$, and the algorithm rejects all values. After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free best-so-far criterion. These two prior-free phases yield robustness. In fact, we derive this design from first principles, as we will explain shortly in Section 1.3.
@@
Suppose $0\leq\beta\leq\frac{1}{e}$ and $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. Further, suppose there is a left-continuous non-increasing function $\theta:[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
$\theta(\lambda_{1})$ $~=~1~,$
$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall\lambda_{1}\leq z\leq\lambda_{2}~.$
Then, there is an $\alpha$-consistent $\beta$-robust algorithm for the MaxExp objective.
To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation in Theorem 1, which does not appear to admit a closed-form solution. See Appendix B.
@@
Suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective for any $\beta\in[0,\frac{1}{e}]$ and
@@
We further prove that certain consistency-robustness trade-offs cannot be achieved for the MaxProb objective. In particular, there is no “best-of-both-world” algorithm.
@@
There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective only if a polytope $\mathcal{P}(\alpha,\beta)$ is feasible. In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\frac{1}{\mathrm{e}}$-robust.
@@ 1.3 Overview of Techniques
A key property of bi-criteria algorithms is that they provide two equivalent viewpoints for robustness. By definition, we may consider a real prior $F$ and a (mis)predicted prior $\tilde{F}$ that misleads the algorithm. Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq(F\circ\tilde{F}^{-1}\circ\theta)(t)$. The latter viewpoint is much easier to work with — we just need to restrict the misspecified threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ via our choice of threshold $\theta$. Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ has the same value of 0 or $1$. This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ is between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any value that is best-so-far. See Section 3 for a detailed discussion, and Section 5 for the robustness analysis.
@@
We show that optimal stopping with a predicted prior is hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest. This is inspired by the sharding technique by harbNewProphetInequalities2025, which breaks each value into many i.i.d. shards based on the prior, such that the original value equals the maximum shard. However, the sharding technique requires full knowledge of the real prior, and thus, does not apply to optimal stopping with a predicted prior. In contrast, implicit sharding directly works with the original values, and only uses the shards implicitly in the analysis through a coupling argument. See Section 4 and specifically Theorem 7 for the formal argument.
@@ 1.4 Related Work
hillComparisonsStopRule1982 proposed a $(1-\frac{1}{\mathrm{e}})$-competitive algorithm for i.i.d. prophet inequality and the MaxExp objective. This was improved to $0.738$ by abolhassaniBeating11Ordered2017, and finally to approximately $0.745$ by correaPostedPriceMechanisms2017. The latter ratio is tight, matching a hardness result by Hill and Kertz [hillComparisonsStopRule1982, kertzStopRuleSupremum1986]. For general prophet inequality with non-identical distributions, the optimal competitive ratio is $0.5$ [krengel1978semiamarts, samuel1984comparison]. The prophet secretary problem [esfandiariProphetSecretary2015] interpolates between general and i.i.d. prophet inequalities, with values drawn from non-identical distributions and arriving by a random order. The optimal competitive ratio for prophet secretary is open, with the best lower and upper bounds being $0.688$ [chenProphetSecretaryMatching2025] and $0.723$ [giambartolomeiProphetInequalitiesSeparating2023].
For the MaxProb objective, gilbertRecognizingMaximumSequence gave an optimal algorithm for the i.i.d. case, while samuels1982 derived an explicit expression of the optimal competitive ratio of approximately $0.580$. For non-identical distributions, esfandiariProphetsSecretariesMaximizing2020 obtained an optimal $\frac{1}{\mathrm{e}}$ competitive ratio in the prophet inequality model, and recovered the $0.580$ ratio in the prophet secretary model under a “no-superstar” assumption. nutiSecretaryProblemDistributions2022 later removed this assumption.
See correaRecentDevelopmentsProphet2019 for a survey. We stress that none of these prophet inequality algorithms provides non-trivial robustness.
@@
There has been extensive research on the sample-based model, in which the algorithm has access to i.i.d. samples from the prior. One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [azar2014prophet, rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022, ezraProphetInequalitySamples2024]. Notably, the optimal $\frac{1}{2}$-competitive prophet inequality can be achieved with only one sample from each distribution [rubinsteinOptimalSinglechoiceProphet2020, gravinOptimalProphetInequality2022]. Another line of work studies the sample complexity of learning a nearly optimal algorithm [diakonikolasLearningOnlineAlgorithms2021, correaUnknownIIDProphets2021, correaTwosidedGameGoogol2020, rubinsteinOptimalSinglechoiceProphet2020, cristiProphetInequalitiesRequire2023, guo2021generalizing, correaSampledrivenOptimalStopping2021, jin2024sample], culminating in the state-of-the-art $\tilde{\mathcal{O}}(\frac{n}{\varepsilon})$ sample complexity for i.i.d. prophet inequality [correaSampledrivenOptimalStopping2021], and $\tilde{\mathcal{O}}(\frac{1}{\varepsilon^{2}})$ sample complexity per distribution for the general case [jin2024sample]. cristiProphetInequalitiesRequire2023 gave a unified proof that $\mathcal{O}(\mathrm{poly}(\frac{1}{\epsilon}))$ samples per distribution suffice for several major variants of prophet inequalities.
Other models of partial information include learning nearly optimal algorithms via quantile queries [liProphetInequalityIID2023], bandit feedback [gatmiry2024bandit], or simple posted-pricing policies [perezsalazarIIDProphetInequality2023]. These models and the sample-based model are incomparable to ours. They assume partial information from the correct prior, while the model of this paper considers a predicted prior that may be arbitrarily wrong.
@@
Our work falls within the algorithms with predictions framework [mitzenmacher2022algorithms] for augmenting worst-case analysis of algorithms with machine-learned advice. Consistency, robustness, and smoothness are the primary performance metrics, capturing the algorithms’ performance when the prediction is correct, completely wrong, and partially wrong. The framework has been applied to many optimization problems, including scheduling [kumar2018improving] and caching [lykouris2021competitive]. In the context of optimal stopping, antoniadisSecretaryOnlineMatching2020 studied a prediction of the maximum value, fujiiSecretaryProblemPredictions2024 assumed having a predicted sequence of values, and braun2024secretary considered a prediction of the gap between the maximum value and the $k$-th largest. In contrast, this paper considers distributional predictions, which aligns with the literature of prophet inequality and the machine-learned priors in practice. As mentioned earlier, duttingPostedPricingProphet2019 also considered mispredicted prior with a focus on smoothness, while this paper examines the trade-off between consistency and robustness.

### v2

@@
Optimal Stopping with a Predicted Prior [^1] This second arXiv version improves the exposition based on feedback we received after posting the first version. We acknowledge the concurrent work by Kehne and Kesselheim [40], add Section 3.3 that explains how to use the indifference Bellman equation to compute the optimal threshold function within our algorithm class, and strengthen Theorem 1 by establishing that the trade‑off curve for MaxProb is tight within the class.
@@ 1 Introduction
Optimal stopping problems study a fundamental dilemma in decision‑making: to seize an expiring opportunity now, or to wait for a better chance that may never come. A classic example is the secretary problem. In this problem, $n$ values are revealed to a decision maker in a random order. After observing each value, the decision maker immediately decides whether to accept the value. The decision maker can accept at most one value and aims to maximize its expectation. An optimal solution is Dynkin’s algorithm [21, 29], which rejects the first $\lfloor\nicefrac{{n}}{{\mathrm{e}}}\rfloor$ values, and then accepts the first value that exceeds all previous ones. It guarantees that the expected accepted value is at least $\nicefrac{{1}}{{\mathrm{e}}}$ times the maximum. This ratio of the algorithm’s expected accepted value to the maximum value in hindsight is called the competitive ratio.
In sharp contrast to the minimal prior knowledge assumed in the secretary problem, prophet inequality considers values drawn independently from fully known prior distributions. In the i.i.d. case where values are drawn from a common distribution, the optimal algorithm guarantees that the expected accepted value is at least approximately $0.745$ times the expected maximum value [38, 41, 16, 44], substantially better than the prior-free secretary guarantee of $\nicefrac{{1}}{{\mathrm{e}}}\approx 0.368$.
On the one hand, the secretary problem does not capture the abundant data available to decision makers in this era. On the other hand, the assumption of complete knowledge of the prior in prophet inequality is idealized. What real-world decision makers often have are machine-learned and potentially erroneous predicted priors. Dütting and Kesselheim [20] took a step in this direction. They show that the performance of the optimal backward-induction algorithm that fully trusts the predicted prior degrades proportionally to some appropriate measures of prediction error, thereby providing a smoothness-style guarantee in the framework of algorithms with predictions [46].
@@ 1.1 Optimal Stopping with a Predicted Prior
The $\nicefrac{{1}}{{\mathrm{e}}}$-competitive Dynkin’s algorithm is $\nicefrac{{1}}{{\mathrm{e}}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust for both objectives. The optimal $\alpha^{*}_{\textrm{MaxExp}}\approx 0.745$-competitive MaxExp algorithm by Hill and Kertz [38, 41] and $\alpha^{*}_{\textrm{MaxProb}}\approx 0.580$-competitive MaxProb algorithm by Gilbert and Mosteller [33] are $\alpha^{*}_{\textrm{MaxExp}}$ and $\alpha^{*}_{\textrm{MaxProb}}$-consistent, respectively, but only 0-robust (see Section 2). Further, randomly running either Dynkin’s algorithm or an i.i.d. prophet inequality algorithm linearly interpolates between their consistency and robustness. Finally, the optimality of these competitive ratios implies that no algorithm is better than $\nicefrac{{1}}{{\mathrm{e}}}$-robust, or better than $\alpha^{*}_{\textrm{MaxExp}}$-consistent for MaxExp, or better than $\alpha^{*}_{\textrm{MaxProb}}$-consistent for MaxProb. These baseline algorithms and impossibilities leave a significant gap between them.
@@ 1.2 Overview of Results
We define a family of bi-criteria algorithms $\mathcal{A}(\tilde{F},\theta)$ parameterized by the predicted prior $\tilde{F}$ and a threshold function $\theta$. Upon observing each value and when no value has been accepted, a bi-criteria algorithm accepts the current value if:
@@
    It is a record, i.e., greater than or equal to all previous values; (prior-free criterion)
@@
For now, readers may interpret $t$ as the time step by which the value is observed. Later, we will also consider an equivalent continuous-time model to simplify the analysis (see Section 2 for a discussion).
We set the thresholds as follows. Before some time $\lambda_{1}$, the threshold is $1$, and the algorithm rejects all values. After some time $\lambda_{2}$, the threshold is 0, and the algorithm’s decisions depend solely on the prior-free record-based criterion. These two prior-free phases yield robustness. In fact, we derive this design from first principles, as we will explain shortly in Section 1.3.
@@
For any $\beta\in[0,\nicefrac{{1}}{{e}}]$, suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. If there is a left-continuous non-increasing function $\theta\colon[\lambda_{1},\lambda_{2}]\to[0,1]$ satisfying:
$\theta(\lambda_{1})$ $~=~1~,$ (1)
$\int_{z}^{1}\int_{0}^{t}\,\frac{1}{t}\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t$ $~\geq~\alpha\cdot\theta(z)~,\qquad\forall z\in[\lambda_{1},\lambda_{2}]~,$
then for any $n\in\mathbb{N}$, there is a bi-criteria algorithm that is $\alpha$-consistent and $\beta$-robust for the MaxExp objective.
The proof of Theorem 1 is given in Section 6. To plot the curve in Figure 1(a), we rely on numerical methods to solve the differential equation (1), which does not appear to admit a closed-form solution. See Appendix B.
@@
For any $\beta\in[0,\frac{1}{e}]$, suppose $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. For any $n\in\mathbb{N}$, there is an $\alpha$-consistent $\beta$-robust bi-criteria algorithm for the MaxProb objective, with
@@
where $\gamma\approx 0.804$ is the solution to $\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!k}=1$. Moreover, this trade-off is tight within the class of bi-criteria algorithms in the sense that, for any specified $(\alpha,\beta)$, none of the bi-criteria algorithms can improve both objectives simultaneously.
The proof of Theorem 2 is given in Section 7.
For the MaxProb objective, besides the tightness within the bi-criteria family above, we further show a non-trivial hardness that certain consistency-robustness trade-offs cannot be achieved by any algorithm. In particular, this implies that there is no “best-of-both-world” algorithm for this objective.
@@
There is an $\alpha$-consistent $\beta$-robust algorithm for the MaxProb objective only if a polytope $\mathcal{P}(\alpha,\beta)$ is feasible. In particular, no algorithm is both $\alpha^{*}_{\textrm{MaxProb}}$-consistent and $\nicefrac{{1}}{{\mathrm{e}}}$-robust.
Again, to plot the curve in Figure 1(b), we rely on numerical linear program solvers to solve the feasibility of certain polytopes, which does not appear to admit a closed-form solution.
@@ 1.3 Overview of Techniques
A key property of bi-criteria algorithms is that they provide two equivalent viewpoints for robustness. By definition, we may consider a real prior $F$ and a (mis)predicted prior $\tilde{F}$ that misleads the algorithm. Equivalently, we may also consider an algorithm with the correct prior $F$ but a misspecified threshold function $F\circ\tilde{F}^{-1}\circ\theta$, since $\tilde{F}(x)\geq\theta(t)$ if and only if $F(x)\geq F\circ\tilde{F}^{-1}\circ\theta(t)$. The latter viewpoint is much easier to work with — we just need to restrict the misspecified threshold $F\circ\tilde{F}^{-1}\circ\theta(t)$ via our choice of threshold $\theta$. Further, the only restrictions are that whenever $\theta(t)$ is 0 or $1$, $F\circ\tilde{F}^{-1}\circ\theta(t)$ should have the same value of 0 or $1$. This observation leads to our design of thresholds with three phases: 1) a prior-free phase before time $\lambda_{1}$ where $\theta(t)=1$ and the algorithm rejects all values; 2) a prior-dependent phase from time $\lambda_{1}$ to $\lambda_{2}$ where $\theta(t)$ lies between 0 and $1$; and 3) a final prior-free phase after time $\lambda_{2}$ where $\theta(t)=0$ and the algorithm takes any record. See Section 3 for a detailed discussion, and Section 5 for the robustness analysis.
@@
We show that optimal stopping with a predicted prior is the hardest when the number of values tends to infinity, using an implicit sharding technique that may be of independent interest. This is inspired by the sharding technique by Harb [37], which breaks each value into many i.i.d. shards based on the prior, such that the original value equals the maximum shard. However, the sharding technique requires full knowledge of the real prior, and thus, does not apply to optimal stopping with a predicted prior. In contrast, implicit sharding directly works with the original values, and only uses the shards implicitly in the analysis through a coupling argument. See Section 4 and specifically Theorem 9 for the formal argument.
@@ 1.4 Related Work
When the algorithm has complete knowledge of the distributions, the prophet inequality problem has been extensively studied. Under the MaxExp objective, Hill and Kertz [38] proposed a $(1-\nicefrac{{1}}{{\mathrm{e}}})$-competitive algorithm for i.i.d. prophet inequality. This was improved to $0.738$ by Abolhassani et al. [1], and finally to approximately $0.745$ by Correa et al. [16]. The latter ratio is tight, matching a hardness result by Hill and Kertz [38, 41]. For prophet inequality where values are independently drawn from non-identical distributions, the optimal competitive ratio is $0.5$ [43, 51]. The prophet secretary problem [23] interpolates between these two settings: values are independently drawn from non-identical distributions and arrive in random order. The optimal competitive ratio for prophet secretary is open: a lower bound of $0.688$ [10] and an upper bound of $0.723$ [32] are currently known.
For the MaxProb objective, Gilbert and Mosteller [33] gave an optimal algorithm for the i.i.d. prophet inequality, while Samuels [52] derived an explicit expression for the optimal competitive ratio of approximately $0.580$. For non-identical distributions, Esfandiari et al. [24] obtained an optimal $\nicefrac{{1}}{{\mathrm{e}}}$-competitive ratio for the prophet inequality model, and recovered the $0.580$ ratio for the prophet secretary under a “no-superstar” assumption. Nuti [47] later removed this assumption.
See Correa et al. [15] for a survey. We stress that none of these prophet inequality algorithms provide non-trivial robustness guarantees.
@@
Relaxing the assumption of complete prior knowledge, several models have studied algorithms with partial information about the distributions. The most prominent one is the sample-based model, in which the algorithm has access to i.i.d. samples from the prior distribution(s). One line of research focuses on getting constant-factor prophet inequalities using one (or fewer) sample from each distribution [4, 50, 35, 26]. Notably, the optimal $\nicefrac{{1}}{{2}}$-competitive prophet inequality can be achieved with only one sample from each distribution [50, 35]. Another line of work studies the minimal number of samples needed to achieve the optimal competitive ratio up to $\varepsilon$ [19, 13, 11, 50, 18, 36, 12, 39]. The state-of-the-art bounds for i.i.d. prophet inequality are $\tilde{\mathcal{O}}(\nicefrac{{n}}{{\varepsilon}})$ for learning a $0.745-\varepsilon$ competitive algorithm [12], and $\tilde{\mathcal{O}}(\nicefrac{{n}}{{\varepsilon^{2}}})$ for learning an online policy that is optimal up to an $\varepsilon$ additive error [36]. For general prophet inequality, $\tilde{\mathcal{O}}(\nicefrac{{1}}{{\varepsilon^{2}}})$ samples per distribution suffice [39]. Finally, Cristi and Ziliotto [18] gave a unified proof that $\mathcal{O}(\mathrm{poly}(\nicefrac{{1}}{{\varepsilon}}))$ samples per distribution suffice for two variants of prophet inequalities. Other models of partial information include quantile-query oracles [28, 48] and bandit feedback [31]. These models and the sample-based model are incomparable to ours: they assume access to the correct prior, while we consider a predicted prior that may be arbitrarily wrong.
@@
Our work falls within the algorithms with predictions framework [46] for augmenting worst-case analysis of algorithms with machine-learned advice. Consistency, robustness, and smoothness are the primary performance metrics, capturing the algorithms’ performance when the prediction is correct, completely wrong, and partially wrong. This framework has been applied to many online optimization problems, including ski-rental [49, 19, 6, 2], scheduling [49, 6], caching [45], online bidding [2], and so on.
Within optimal stopping, prior work has explored several forms of predictions. Antoniadis et al. [3] studies a prediction of the maximum value, Fujii and Yoshida [30] and Balkanski et al. [5] assume having a predicted sequence of values, and Braun and Sarkar [7] considers a prediction of the gap between the maximum and the $k$-th largest value.
In contrast, our work considers distributional predictions, bridging the extensive literature of prophet inequality with the growing use of machine-learned priors in practice. Within the context of prophet inequalities, the only relevant work to our knowledge is by Dütting and Kesselheim [20], which examines the impact of mispredicted priors with an emphasis on smoothness. This work is the first to investigate the consistency-robustness trade-off given a distributional prediction. In general, while distributional predictions have been explored in certain domains such as online matching, where they have been incorporated into traditional stochastic models [25, 9], this type of prediction remains largely underexplored.
Independent and Concurrent Work.
Kehne and Kesselheim [40] study the same problem, focusing on the MaxExp setting. Their results and ours are complementary. The algorithms in both papers are under the bi-criteria framework. They show a non-trivial consistency-robustness trade-off using a simple single-threshold between $\lambda_{1}$ and $\lambda_{2}$, whereas we optimize the choice of thresholds within the bi-criteria framework. They rule out “best-of-both-world” algorithms for the MaxExp setting, whereas we prove such an impossibility result for the MaxProb setting.